if the ratio of 2 complementary angles is 8:1, what is the degree measure of the samller angle
Answer by stanbon(75887)

if the ratio of 2 complementary angles is 8:1, what is the degree measure of the samller angle
---------------------------
Let one of the ansgles be "x"; the other is "90-x".
----
Equation:
x/(90-x) = 8/1
x = 8(90-x)
---
x = 720-8x
9x = 720
x = 80
---
90-x = 10 degrees (smaller angle)
======================================
Cheers,
Stan H.